Sue heads due west from Dry Gulch on U.S. 6.5 at a steady speed of 55 mph (miles per hour) while Don heads due east from Wet Hollow (which is 830 miles west of Dry Gulch) on the same hiway at 65 mph but he starts 2 hours after Sue did. How many miles did Don travel before they met?
Answer by khwang(438)

Two hours after Sue started, their distance is 850 - 55*2 = 740 miles.
Then , another T = 740/(55+65) = 740/120 = 37/6 hrs, they met.
And so Don drove 65 * 37/6 = 2405/6 miles before thy met.

Kenny